the numbers of seats in the first 12 row of a high school auditorium form an aritmetic sequence the first row has 9 seats the second row has 11 seats write a recursive formula to represent the sequence. write an expicit formula to represent the sequence. how many seats are in the 12th row
Answer by Herdaysoji(25)

first term of an A.P (a)=9
second term, (T2)=a+d,11=9+d
d=11-9=2
the 12th term, (T12)=a+(n-1)d
.'.(T12)=a+(12-1)d, =9+(11).2=31
the 12th ter mm is 31